OpenStudy (anonymous):

Help please! solve the definite integral integral of e^(2x)/(1+e^(2x)) dx from 2 to 4

OpenStudy (luigi0210):

Do a u-sub.

OpenStudy (anonymous):

so the u value would be 1+e^(2x)?

OpenStudy (luigi0210):

Yes, now find what du would be.

OpenStudy (anonymous):

so du=2(e^(2x)) dx

OpenStudy (anonymous):

which then becomes 1/2du=e^(2x)dx

OpenStudy (luigi0210):

Right! So now all you have to integrate is \[\LARGE \frac{1}{2} \int \frac{1}{u}~du\]

OpenStudy (luigi0210):

Forgot the limits \[\LARGE \frac{1}{2} \int_{2}^{4} \frac{1}{u}~du\]

OpenStudy (anonymous):

so that would be... \[1/2 \int\limits_{2}^{4}\ln u\]

OpenStudy (anonymous):

oops sorry i mean \[(1/2)\ln(u)\]

OpenStudy (luigi0210):

Yup, now plug in what u is and evaluate it.. \[\LARGE \frac{ln|1+e^{2x}|}{2}|_{2}^{4}\]

OpenStudy (anonymous):

so the answer would be...about 2

OpenStudy (luigi0210):

1.9910927 to be exact, but approx 2 yes :)

OpenStudy (anonymous):

awesome :D would leaving it just at 2 be ok?

OpenStudy (luigi0210):

Depends on what your instructions or instructor wants.. USUALLY they like rounded decimal answers. But just 2 should be fine.

OpenStudy (anonymous):

ok awesome :) thank you so much!

OpenStudy (luigi0210):

You're welcome, and good luck! :P